Luisa mows lawns during the summer. She charges $15 if she cuts the grass but charges $5 more if she also trims the grass. Last week she trimmed 5 more yards than she cut,. If she made $415 last week, how may yards did she trim?
Answer by nerdybill(7384)

Luisa mows lawns during the summer. She charges $15 if she cuts the grass but charges $5 more if she also trims the grass. Last week she trimmed 5 more yards than she cut,. If she made $415 last week, how may yards did she trim?
.
Let x = number of yards just cut
then
x+5 = number of yards cut and trimmed
.
15x + 20(x+5) = 415
15x + 20x+100 = 415
35x+100 = 415
35x = 315
x = 315/35
x = 9 (just cut)
.
Yards trimmed and cut:
x+5 = 9+5 = 14